OpenStudy (anonymous):

find the derivative of y=ln[(√(sinx cosx@)/(1+2lnx)]

OpenStudy (anonymous):

You can simplify the process by writing the problem as ln(sqrt(sinxcosx)) - ln(1 + 2lnx); now the first term is ln(sinxcosx)^1/2 = (1/2)ln(sinxcosx) = 1/2 (lnsinx + ln cosx)

OpenStudy (anonymous):

good luck @emailsolah2

OpenStudy (anonymous):

tq. got it :D